How many flourine molecules are required to produce 2mg of uf6 from an exess of uranium ? the molar mass of uf6 is 352 g mol^-1
eq..
u+3f2_uf6

Grade:11

1 Answers

Suraj Prasad IIT Patna
7 years ago
U + 3F2 = UF6

I have to produce 2*10-3g ofUF6

So no. of moles of UF6 to produce is 2*10^-3 / 352 = 5*10^-6.

3*NAmolecules of fluorine gives 1 mole of UF6

this implies 5*10^-6 moles of UF6 will require 3*6.022*10^23* 5* 10^-6 molecules of fluorine which comes out to be 9.033*10^18 molecules of fluorine.